how to solve 2y(y+6)=32

Answer by stanbon(75887)

how to solve 2y(y+6)=32
--
y(y+6) = 16
y^2+6y-16 = 0
-----
Factor:
(y+8)(y-2) = 0
y = -8 or y = 2
========================
Cheers,
Stan H.